OpenStudy (egracer):

Consider the functions f(x) = 3x2 and g(x) = −2x − 5. What is the value of f[g(−4)]? 9 −31 27 3

OpenStudy (anonymous):

calculate g(-4) first

OpenStudy (egracer):

How?

OpenStudy (anonymous):

g(-4) = -2 (-4) -5 = 8 - 5 = 3

OpenStudy (anonymous):

now substitute 3 in f(x)

OpenStudy (anonymous):

f(3) = 3* 3^2 = 27

OpenStudy (anonymous):

so that's the answer

OpenStudy (egracer):

thank you so much! I think I understand the whole thing now! hopefully ill be able to do the next one on my own haha

OpenStudy (anonymous):

Good luck with them!!